### v1

# TwinFlow: Realizing One-step Generation on Large Models with Self-adversarial Flows

Abstract
Recent advances in large multi-modal generative models have demonstrated impressive capabilities in multi-modal generation, including image and video generation. These models are typically built upon multi-step frameworks like diffusion and flow matching, which inherently limits their inference efficiency (requiring 40-100 Number of Function Evaluations (NFEs)). While various few-step methods aim to accelerate the inference, existing solutions have clear limitations. Prominent distillation-based methods, such as progressive and consistency distillation, either require an iterative distillation procedure or show significant degradation at very few steps (< 4-NFE). Meanwhile, integrating adversarial training into distillation (e.g., DMD/DMD2 and SANA-Sprint) to enhance performance introduces training instability, added complexity, and high GPU memory overhead due to the auxiliary trained models. To this end, we propose TwinFlow, a simple yet effective framework for training 1-step generative models that bypasses the need of fixed pretrained teacher models and avoids standard adversarial networks during training, making it ideal for building large-scale, efficient models. On text-to-image tasks, our method achieves a GenEval score of 0.83 in 1-NFE, outperforming strong baselines like SANA-Sprint (a GAN loss-based framework) and RCGM (a consistency-based framework). Notably, we demonstrate the scalability of TwinFlow by full-parameter training on Qwen-Image-20B and transform it into an efficient few-step generator. With just 1-NFE, our approach matches the performance of the original 100-NFE model on both the GenEval and DPG-Bench benchmarks, reducing computational cost by $100\times$ with minor quality degradation. Project page is available here.
Code: https://github.com/inclusionAI/TwinFlow
[Uncaptioned image]Models: https://huggingface.co/inclusionAI/TwinFlow
Figure 1: Results of Qwen-Image-20B-TwinFlow (NFE=2). See prompts in App. D.3.
1 Introduction
Modern generative paradigms—including diffusion (Ho et al., 2020; Song et al., 2020a), flow matching (Lipman et al., 2022; Ma et al., 2024), and consistency models (Song et al., 2023; Lu & Song, 2024)—have achieved state-of-the-art performance, forming the backbone of leading image and video generation systems (Peebles & Xie, 2023; Ho et al., 2022; Chen et al., 2024c; Xie et al., 2024a). Despite their success, these methods share a critical drawback: both training and sampling demand substantial computational resources. This challenge is magnified in the era of large-scale models (ModelTC, 2025). For these systems, efficient sampling is paramount, as the continuous, long-term cost of inference often surpasses the one-time training cost, directly impacting their economic viability and practical deployment (ModelTC, 2025; Xie et al., 2025a).
Numerous research efforts aim to accelerate generative inference by reducing the number of sampling steps. Early single-step methods, like Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), often suffered from unstable training dynamics. To accelerate the multi-step diffusion models, various distillation techniques have been introduced. These range from student-teacher methods, where a compact model learns to emulate a larger one in fewer steps (Salimans & Ho, 2022; Meng et al., 2022), to distribution matching distillation (e.g., DMD variants (Yin et al., 2024b; a)), which uses adversarial training to directly align the model’s output distribution with the real data. In parallel, a powerful new paradigm of consistency models (Song et al., 2023) and their variants, such as LCM (Luo et al., 2023) and PCM (Wang et al., 2024a), has emerged, explicitly designed for high-quality generation in very few steps.
Despite their progress, existing few-step methods face a difficult trade-off between simplicity, efficiency, and quality. (a) Complexity and instability: As summarized in Tab. 1, adversarial methods such as GANs and DMD require auxiliary networks (e.g., discriminators) or frozen teacher models. This not only introduces training instability and sensitivity to hyperparameters but also increases architectural complexity and memory overhead (c.f. Fig. 2(b)), hindering their scalability to large models. (b) Performance degradation: Conversely, methods that train from scratch without adversarial guidance (Luo et al., 2023; Yin et al., 2024a), such as consistency models, often exhibit a sharp decline in quality at very low NFEs (< 4) (Chen et al., 2025b). In summary, we posit that existing methods either suffer from training instability, or require additional/frozen models (see our Tab. 1), which limits their simplicity and scalability in training large models.
Table 1: Comparison of different few-step generative modeling methods on their minimal dependence of auxiliary trained model and frozen teacher model. Prior 1-step/few-step methods such as GAN requires a trained discriminator, diffusion/consistency distillation requires a frozen teacher model, DMD requires training an auxiliary score function for fake data and a frozen teacher model, DMD2 trains a GAN discriminator and a fake score function at the same time. Our TwinFlow achieves 1-step generation without depending on auxiliary trained or frozen models, offering high simplicity.
Method Generation type #Auxiliary #Frozen
trained model teacher model
GAN (Goodfellow et al., 2014) 1-step 1 0
Diffusion models (Ho et al., 2020) multi-step 0 0
Flow matching models (Lipman et al., 2022) multi-step 0 0
Diffusion distillation (Salimans & Ho, 2022) few-step 0 1
Consistency training & distillation (Song et al., 2023) 1-step, few-step 0 0,1
Distribution matching distillation (Yin et al., 2024b; a) 1-step, few-step 1,2 1
TwinFlow (Ours) 1-step, few-step 0 0
To address these challenges, we propose TwinFlow, a simple yet effective one-step generative training framework built on a novel twin-trajectory concept. By extending the standard time interval from $t\in[0,1]$ to $t\in[-1,1]$, we conceptualize two trajectories originating from the noise distribution. The positive branch ($t>0$) maps noise to real data while the negative branch ($t<0$) maps the same noise to “fake” data, enabling simultaneous learning of both transformations. Our learning objective is to minimize the discrepancy between the velocity fields of these two trajectories (see Fig. 2(a)). This forces the model to learn a more robust and direct mapping from noise to data, thereby enhancing 1-step generation performance in a self-supervised manner. As highlighted in Tab. 1, a key advantage of TwinFlow is its simplicity, as it requires no auxiliary trained networks or frozen teacher models. Extensive experiments at different scales demonstrate the effectiveness of TwinFlow, including text-to-image generation (cf. Sec. 4.2 & Sec. 4.3) on large models like Qwen-Image-20B (Wu et al., 2025a). 2-NFE visualizations on Qwen-Image-20B are given in Fig. 1. Our key contributions are:
  1. (a)
    Simple yet effective 1-step generation framework. We propose a one-step generation framework that does not need auxiliary trained models (GAN discriminators) or frozen teacher models (different/consistency distillation), thereby eliminating GPU memory cost, allowing for more flexible and scalable training on large models.
  2. (b)
    Strong 1-NFE performance on text-to-image task. Built on the any-step framework, TwinFlow achieves strong text-to-image performance with only 1-NFE, achieving 0.83 GenEval score, surpassing SANA-Sprint (0.72) and RCGM (0.80).
  3. (c)
    Effective application on large models. By applying TwinFlow, we successfully bring 1/2-NFE generation capabilities to Qwen-Image-20B, the largest open-source multi-modal generation model. We achieve GenEval score of 0.86 and DPG score of 86.52 (1-NFE); GenEval 0.87 and DPG Score 87.64 (2-NFE), which are highly competitive with the original 100-NFE scores of 0.87 and 88.32.
(a) TwinFlow overview. The standard flow is on the right side (solid lines), its twin (dashed lines) is on the left side. The core of our method is to minimize of the difference between the velocity fields ($\Delta_{\mathbf{v}}$) of the standard flow and its twin flow.
(b) GPU memory comparison. Directly adopting DMD2 and SANA-Sprint suffers from OOM when applying to ultra-large models. Our method can be easily applied to train Qwen-Image-20B.
Figure 2: Overview of our TwinFlow and training GPU memory comparison. The GPU memory usage is measured on $\times1024$ resolution on Qwen-Image-20B (LoRA tuning) and SANA-1.6B.
2 Preliminaries
Given a dataset $\mathcal{D}$, let $p(\mathbf{x})$ represent its data distribution and $p(\mathbf{x}|\mathbf{c})$ the conditional distribution given a condition $\mathbf{c}$. Generative models aim to learn a transformation from a simple source distribution, $p(\mathbf{z})$, such as the standard Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, to the complex target distribution, $p(\mathbf{x})$.
Any-step generative model framework.
A recent framework, RCGM (Sun & Lin, 2025), introduces a unified formulation for the any-step generation framework, covering paradigms like multi-step generative models (Ho et al., 2020; Song et al., 2020b; Lipman et al., 2022) and few-step generative models (Song et al., 2023; Lu & Song, 2024; Frans et al., 2024; Geng et al., 2025; Sun et al., 2025).
In this framework, a prediction function can be generally defined as $\boldsymbol{f}(\mathbf{x}_{t},r):=\mathbf{x}_{r}-\mathbf{x}_{t}$, which predicts the target point $\mathbf{x}_{r}$ from the current point $\mathbf{x}_{t}$ along a specific PF-ODE trajectory. The unified training objective for any-step models is given by:
$\mathcal{L}(\boldsymbol{\theta})_{\mathrm{base}}={\mathbb{E}}_{\mathbf{x},\mathbf{z},\{t_{i}\}_{i=0}^{N+1}}\left[d\left(\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t},\frac{1}{\Delta t}\left[\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t_{N+1})-\sum_{i=1}^{N}\boldsymbol{f}_{\boldsymbol{\theta}^{-}}(\mathbf{x}_{t_{i}},t_{i+1})\right]\right)\right]\,$, (1)
where $\mathbf{x}_{t}=\alpha(t)\mathbf{z}+\gamma(t)\mathbf{x}$, $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $t\sim\mathrm{U}(0,T)$, $t_{i}\sim\mathrm{U}(t_{i-1},0)$, and $d(\cdot,\cdot)$ is a metric function. Under flow matching objective and linear transport, we have $\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{t},r)=\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},r)\cdot(t-r)$, where $\boldsymbol{F}_{\boldsymbol{\theta}}$ is a neural network, $\boldsymbol{F}_{\boldsymbol{\theta}^{-}}$ is the no grad version. In practice, we use Network(x_t, t, r) as the implementation of $\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},r)$. Equation (1) demonstrates how both multi-step and few-step frameworks can be seen as specific instances of the broader any-step framework, which we will detail below.
Multi-step generative models.
Diffusion (Ho et al., 2020; Song et al., 2020b) and flow matching models (Lipman et al., 2022) can be derived from the RCGM framework. By setting $N=0$, the objective in (1) reduces to their respective training objectives, where the predict function becomes $\boldsymbol{f}(\mathbf{x}_{t},t-\Delta t)$ in the limit $\Delta t\to 0$. During sampling, these models iteratively solve the PF-ODE by integrating the velocity field $\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$, starting from a noise sample $\mathbf{x}_{1}\sim p(\mathbf{z})$ at $t=1$ and ending at $t=0$ to obtain samples from $p(\mathbf{x})$.
Few-step generative models.
Few-step models are also instances of the RCGM framework, typically corresponding to $N=1$ case: (1) setting $t_{1}=t-\Delta t,\Delta t\to 0$ and $t_{2}=0$ recovers the objective for consistency models (Song et al., 2023; Lu & Song, 2024); (2) setting $t_{1}\in(t_{2},t)$ corresponds to shortcut models (Frans et al., 2024), where the predict function is defined as $\boldsymbol{f}(\mathbf{x}_{t},r)\leftarrow\boldsymbol{f}(\mathbf{x}_{t},s)+\boldsymbol{f}(\mathbf{x}_{r},s)$ and $r=t_{2}\in[0,t]$; (3) setting $t_{1}=t-\Delta t$ (with $\Delta t\to 0$) yields the MeanFlow objective (Geng et al., 2025).
In summary, the RCGM framework offers a unified perspective that integrates both multi-step and few-step paradigms, facilitating their analysis and application. See more related work discussion in Sec. A.
3 Methodology
Current few-step methods within the any-step framework (Sec. 2) struggle to achieve high-quality one-step generation without resorting to a GAN loss, which adds significant complexity. To solve this, we propose TwinFlow, a simple and self-contained approach that enhances one-step performance directly within the any-step flow matching framework. Our key idea is the introduction of twin trajectories, which create an internal self-adversarial signal and thus eliminate the need for an external GAN loss (Sec. 3.1). The method works by minimizing a difference between a “fake” and a “real” velocity field, which should ideally be zero (Sec. 3.2). We conclude by demonstrating how to integrate TwinFlow into the broader any-step framework and provide practical designs in Sec. 3.3.
3.1 Twin Trajectory for Self-adversarial Training
A key innovation of our method is the introduction of twin trajectories, which feature time-steps symmetric around $t=0$ (see Fig. 2(a)). This structure creates a self-contained, discriminator-free adversarial objective designed to directly enhance one-step generation performance.
Creating self-adversarial objective.
The standard learning process operates on a time interval $t\in[0,1]$: real data $\mathbf{x}$ is perturbed by $\mathbf{x}_{t}^{\mathrm{real}}=\alpha(t)\mathbf{z}+\gamma(t)\mathbf{x}$, where $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, $t\sim\mathrm{U}(0,1)$. To create our self-adversarial objective (as well as the twin trajectories), we extend this time interval from $t\in[0,1]$ to $t\in[-1,1]$. The negative half of this interval, $t\in[-1,0]$, is designated for learning a generative path from noise to “fake” data produced by the model itself.
Specifically, we task the network to learn the generative path to its own outputs. We take a fake sample $\mathbf{x}^{\mathrm{fake}}$ generated by the model, i.e. $\mathbf{x}^{\mathrm{fake}}=\hat{\mathbf{x}_{t}}=\boldsymbol{F}_{\theta}(\mathbf{x}_{t}^{\mathrm{real}},t)$, and construct a corresponding “fake trajectory”, in which its perturbed version is defined as $\mathbf{x}^{\mathrm{fake}}_{t^{\prime}}=\alpha(t^{\prime})\mathbf{z}^{\mathrm{fake}}+\gamma(t^{\prime})\mathbf{x}^{\mathrm{fake}}$, $\mathbf{z}^{\mathrm{fake}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $t^{\prime}\sim\mathrm{U}(0,1)$. Here $\mathbf{z}^{\mathrm{fake}}$ is a different noise, which does not need to be the same as $\mathbf{z}$. The network is then trained with the following flow matching objective on this trajectory, using negative time inputs $t^{\prime}\in[-1,0]$:
$\mathcal{L}(\boldsymbol{\theta})_{\mathrm{adv}}={\mathbb{E}}_{\mathbf{x}^{\mathrm{fake}},\mathbf{z}^{\mathrm{fake}},t^{\prime}}\left[d\left(\boldsymbol{F}_{\theta}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}},-t^{\prime}),\mathbf{z}^{\mathrm{fake}}-\mathbf{x}^{\mathrm{fake}}\right)\right]\,$, (2)
where $d(\cdot,\cdot)$ is a metric function. Minimizing this loss teaches the network to learn the negative time condition and the transformation from noise to fake data distribution, setting the stage for the rectification loss described in the next section.
3.2 Rectifying Real Trajectory via Velocity Matching
Ideally, we want the twin trajectories to match with each other. As established in Sec. 3.1, the distributions $p_{\mathrm{fake}}$ and $p_{\mathrm{real}}$ correspond to trajectories parameterized by the negative and positive time intervals, respectively. Inspired by DMD (Yin et al., 2024b), we can treat this as a distribution matching problem. For any perturbed sample $\mathbf{x}_{t}$, we aim to minimize the KL divergence between these two distributions:
$\textstyle D_{\mathrm{KL}}(p_{\mathrm{fake}}\|p_{\mathrm{real}})={\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[\log{\left(\frac{p_{\mathrm{fake}}(\mathbf{x}_{t})}{p_{\mathrm{real}}(\mathbf{x}_{t})}\right)}\right]$ $={\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[-\left(\log{p_{\mathrm{real}}(\mathbf{x}_{t})}-\log{p_{\mathrm{fake}}(\mathbf{x}_{t})}\right)\right]\,.$ (3)
Velocity matching as distribution matching.
Taking the gradient of (3), we derive:
$\textstyle\nabla_{\boldsymbol{\theta}}D_{\mathrm{KL}}(p_{\mathrm{fake}}\|p_{\mathrm{real}})$ $=\nabla_{\boldsymbol{\theta}}{\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[\log p_{\mathrm{fake}}(\mathbf{x}_{t})-\log p_{\mathrm{real}}(\mathbf{x}_{t})\right]$
$=\nabla_{\boldsymbol{\theta}}{\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[\frac{\partial\log p_{\mathrm{fake}}(\mathbf{x}_{t})}{\partial\mathbf{x}_{t}}\;\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}-\frac{\partial\log p_{\mathrm{real}}(\mathbf{x}_{t})}{\partial\mathbf{x}_{t}}\;\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}\right]$
$=\nabla_{\boldsymbol{\theta}}{\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[\Big(\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{\mathrm{fake}}(\mathbf{x}_{t})-\nabla_{\mathbf{x}_{t}}\log p_{\mathrm{real}}(\mathbf{x}_{t})}_{\displaystyle\mathbf{s}_{\mathrm{fake}}(\mathbf{x}_{t})-\mathbf{s}_{\mathrm{real}}(\mathbf{x}_{t})}\Big)\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}\right]\,$, (4)
where $\mathbf{s}(\cdot)$ is the score of the respective distribution. The relationship between the score and the velocity field $\boldsymbol{F}_{\boldsymbol{\theta}}$ under linear transport ($\alpha(t)=t,\gamma(t)=1-t$) is given by (see proof in App. C.1):
$\mathbf{s}(\mathbf{x}_{t})=-\frac{\mathbf{x}_{t}+\gamma(t)\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)}{\alpha(t)}=-\frac{\mathbf{x}_{t}+(1-t)\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)}{t}\,.$ (5)
Substituting this relationship from (5) into the KL gradient (4) yields:
$\textstyle\nabla_{\boldsymbol{\theta}}D_{\mathrm{KL}}$ $=\nabla_{\boldsymbol{\theta}}{\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[\left(-\frac{\mathbf{x}_{t}+(1-t)\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-t})}{t}-(-\frac{\mathbf{x}_{t}+(1-t)\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}t})}{t})\right)\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}\right]$
$=\nabla_{\boldsymbol{\theta}}{\mathbb{E}}_{\mathbf{x}_{t},\mathbf{z},t}\left[-\frac{(1-t)}{t}\cdot\Big(\underbrace{\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-t})}_{\displaystyle\mathbf{v}_{\mathrm{fake}}(\mathbf{x}_{t},-t)}-\underbrace{\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}t})}_{\displaystyle\mathbf{v}_{\mathrm{real}}(\mathbf{x}_{t},t)}\Big)\frac{\partial\mathbf{x}_{t}}{\partial\boldsymbol{\theta}}\right]\,$, (6)
where the model is conditioned on $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-t$} for the fake trajectory and on $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}t$} for the real one. For simplicity, we denote this velocity difference (see Fig. 2(a)) as:
$\Delta_{\mathbf{v}}(\mathbf{x}_{t}):=\mathbf{v}_{\mathrm{fake}}(\mathbf{x}_{t},-t)-\mathbf{v}_{\mathrm{real}}(\mathbf{x}_{t},t)\,.$ (7)
This derivation recasts the original distribution matching problem into a more practical velocity matching problem. We now show how to formulate this into a tractable rectification loss below.
Rectification loss derivation.
To derive the rectification loss, we first instantiate the gradient (6) using the setup in Sec. 3.1. In this setting, the network’s prediction $\hat{\mathbf{x}}_{t}$ serves as the clean example, and consequently, the perturbed variable $\mathbf{x}_{t}$ in (6) corresponds to the fake sample $\mathbf{x}^{\mathrm{fake}}_{t^{\prime}}$. The velocity difference $\Delta_{\mathbf{v}}(\mathbf{x}_{t})$ defined in (7) is therefore instantiated as $\Delta_{\mathbf{v}}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}})=\mathbf{v}_{\mathrm{fake}}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}},-t^{\prime})-\mathbf{v}_{\mathrm{real}}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}},t^{\prime})$.
Under this setup, the Jacobian term in (6) is instantiated as $\frac{\partial\mathbf{x}^{\mathrm{fake}}_{t^{\prime}}}{\partial\boldsymbol{\theta}}$ and simplified to:
$\frac{\partial\mathbf{x}^{\mathrm{fake}}_{t^{\prime}}}{\partial\boldsymbol{\theta}}=\frac{\partial(\alpha(t^{\prime})\mathbf{z}^{\mathrm{fake}}+\gamma(t^{\prime})\mathbf{x}^{\mathrm{fake}})}{\partial\boldsymbol{\theta}}=\frac{\partial(\alpha(t^{\prime})\mathbf{z}^{\mathrm{fake}}+\gamma(t^{\prime})\hat{\mathbf{x}_{t}})}{\partial\boldsymbol{\theta}}=\frac{\partial\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t}^{\mathrm{real}},t)}{\partial\boldsymbol{\theta}}\,.$ (8)
The KL gradient in (6) thus takes the form of an expectation over the inner product $\langle\Delta_{\mathbf{v}}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}}),\frac{\partial\boldsymbol{F}_{\boldsymbol{\theta}}}{\partial\boldsymbol{\theta}}\rangle$. To construct a tractable loss that produces this gradient structure, we employ the stop-gradient operator, $\mathrm{sg}(\cdot)$. This motivates the following rectification loss:
$\mathcal{L}(\boldsymbol{\theta})_{\mathrm{rectify}}={\mathbb{E}}_{\mathbf{x}_{t},\mathbf{x}^{\mathrm{fake}},\mathbf{z}^{\mathrm{fake}},t^{\prime}}\left[d\left(\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t}^{\mathrm{real}},t),\mathrm{sg}\left[\Delta_{\mathbf{v}}(\mathbf{x}^{\mathrm{fake}}_{t^{\prime}})+\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t}^{\mathrm{real}},t)\right]\right)\right]\,$, (9)
where $d(\cdot,\cdot)$ is a metric function. Minimizing $\mathcal{L}_{\mathrm{rectify}}$ encourages the model to straighten the generative trajectories from noise to the data distribution. This rectification allows the entire integration process to be accurately approximated with large step sizes, enabling few-step or 1-step generation.
3.3 The TwinFlow Objective with Practical Designs
Integration with the any-step framework.
Our method TwinFlow trains a single model to excel at both multi-step and few-step generation. This is achieved by combining two complementary objectives with conflicting demands:
  • The self-adversarial loss ($\mathcal{L}(\boldsymbol{\theta})_{\mathrm{adv}}$ in (2)) promotes high-fidelity, multi-step generation by extending the training dynamics to the interval $t\in[-1,0]$.
  • The rectification loss ($\mathcal{L}(\boldsymbol{\theta})_{\mathrm{rectify}}$ in (9)) optimizes for few-step efficiency by directly straightening the noise-to-data trajectory, enabling rapid, high-quality synthesis.
This creates a dual objective: the model must be both a precise multi-step sampler and an efficient few-step generator. This leads to application of the any-step framework introduced in Sec. 2, which unifies the demands of (2) and (9). We adopt N=2N\!=\!2 formulation of (1) to enhance the training stability.
Our final loss combines the base objective with our proposed terms, which we collectively name it $\mathcal{L}(\boldsymbol{\theta})_{\mathrm{TwinFlow}}$. The overall loss function in our methodology can be expressed as:
$\mathcal{L}(\boldsymbol{\theta})=\mathcal{L}(\boldsymbol{\theta})_{\mathrm{base}}+\left(\mathcal{L}(\boldsymbol{\theta})_{\mathrm{adv}}+\mathcal{L}(\boldsymbol{\theta})_{\mathrm{rectify}}\right)=\mathcal{L}(\boldsymbol{\theta})_{\mathrm{base}}+\mathcal{L}(\boldsymbol{\theta})_{\mathrm{TwinFlow}}\,.$ (10)
Practical implementation of mixed loss.
The $\mathcal{L}(\boldsymbol{\theta})_{\mathrm{base}}$ and $\mathcal{L}(\boldsymbol{\theta})_{\mathrm{TwinFlow}}$ objectives in $\mathcal{L}(\boldsymbol{\theta})$ impose different requirements on the target time $r$ under the any-step formulation. Specifically, $\mathcal{L}(\boldsymbol{\theta})_{\mathrm{base}}$ requires $r$ to be sampled from $[0,1]$, whereas $\mathcal{L}(\boldsymbol{\theta})_{\mathrm{TwinFlow}}$ necessitates a fixed target time of $r=0$. To accommodate both within a single training step, we partition each mini-batch into two subsets.
A balancing hyperparameter $\lambda$ controls the relative size of these subsets. One portion of the batch is used to compute $\mathcal{L}(\boldsymbol{\theta})_{\mathrm{TwinFlow}}$ with $r=0$, while the remainder is used for $\mathcal{L}(\boldsymbol{\theta})_{\mathrm{base}}$ with a randomly sampled $r\in[0,1]$. The value of $\lambda$ thus balances the influence of the two losses on the gradient updates. Setting $\lambda=0$ disables the $\mathcal{L}(\boldsymbol{\theta})_{\mathrm{TwinFlow}}$ term, while larger values increase its contribution. An ablation study on the impact of $\lambda$ is available in Fig. 4(a).
4 Experiments
We demonstrate the effectiveness of our method, TwinFlow, on two fronts. First, we highlight its versatility and scalability, we apply TwinFlow to unified multi-modal models, e.g. Qwen-Image-20B (Wu et al., 2025a), as shown in Tab. 2. Second, we benchmark it against state-of-the-art (SOTA) dedicated text-to-image models, with results presented in Tab. 4.
4.1 Experimental Setup
This section details the experimental setup and evaluation protocol of our proposed methodology.
  • Image generation on multimodal generative models. We conduct evaluations on unified multi-modal models (i.e. takes both texts and images as conditions and capable of generating texts and images). (1) Network architectures: We conducted LoRA (Hu et al., 2022) (Tab. 2) and full-parameter training (Tab. 3) of TwinFlow on Qwen-Image. We also do full-parameter training experiments on OpenUni-512 (Wu et al., 2025c). (2) Benchmarks: Following recent works (Pan et al., 2025; Chen et al., 2025a; Deng et al., 2025; Wu et al., 2025a), we use benchmarks in text-to-image generation tasks. For text-to-image generation, we use GenEval, DPG-Bench (Hu et al., 2024), and WISE (Niu et al., 2025). Other training settings are detailed in App. B.1.
  • Text-to-image generation. For text-to-image generation, we evaluate on dedicated text-to-image models (i.e. primarily takes texts as condition and only generating images). (1) Network architectures: We use SANA-0.6B/1.6B (Xie et al., 2024a) in our experiments. (2) Benchmarks: Following SANA-series (Xie et al., 2024a; 2025a), we use GenEval (Ghosh et al., 2023) and DPG-Bench (Hu et al., 2024) as evaluation metrics. Other training settings are detailed in App. B.2.
4.2 Image Generation on Multimodal Generative Models
We demonstrate TwinFlow’s scalability by achieving competitive 1-NFE text-to-image generation on the 20B-parameter Qwen-Image series (Wu et al., 2025a). This breakthrough addresses a critical gap in the field, as prior few-step approaches are rarely applied on models exceeding 3B parameters due to instability in GAN-based loss at scale.
Our approach offers two key advantages over state-of-the-art unified multimodal generative models:
  1. (a)
    TwinFlow maintains >0.86 GenEval score at 1-NFE on Qwen-Image-20B: surpassing most multi-step models (w/ 40-100 NFEs), e.g. Bagel (Deng et al., 2025), MetaQuery (Pan et al., 2025).
  2. (b)
    TwinFlow achieves this without auxiliary components or architectural modifications, unlike competing few-step methods that require distillation or specialized training pipelines (Yin et al., 2024b; a).
We evaluate the text-to-image generation capabilities of Qwen-Image-TwinFlow on several standard benchmarks: GenEval (Ghosh et al., 2023), DPG-Bench (Hu et al., 2024), and WISE (Niu et al., 2025). Our model demonstrates strong performance across all benchmarks with only 1-NFE, achieving results that are both competitive and efficient. Detailed results are provided in App. B.1.
Evaluation on text-to-image benchmarks.
As shown in Tab. 2, Qwen-Image-TwinFlow achieves a score of 0.86 on GenEval and 86.52% on DPG-Bench with just 1-NFE, closely matching the original model’s performance at 100-NFE. Compared to Qwen-Image-Lightning (ModelTC, 2025), a 4-step distilled model, our model surpasses it on GenEval and WISE with only 1-NFE. Furthermore, our model outperforms Qwen-Image-RCGM (Sun & Lin, 2025) on both GenEval and DPG-Bench under 1-NFE and 2-NFE settings, with notable improvements of 0.34 on GenEval, 27.0% on DPG-Bench, and 0.25 on WISE under the 1-NFE setting.
We also benchmark Qwen-Image-TwinFlow against other prominent multi-step unified multimodal generative models, such as MetaQuery-XL (Pan et al., 2025), BLIP3-o-8B (Chen et al., 2025a), and Bagel (Deng et al., 2025). Our model consistently surpasses these baselines with 1 or 2-NFE across all evaluation metrics. Beyond Qwen-Image, we also apply TwinFlow to OpenUni (Wu et al., 2025c), achieving GenEval of 0.80 and DPG-Bench of 76.40 under the 1-NFE setting, which is also close to its original performance. These findings underscore the versatility and effectiveness of TwinFlow across different architectures and scales.
Table 2: System-level comparison of TwinFlow with unified multimodal models in efficiency and performance on text-to-image tasks. The best and second best results of 1-NFE and 2-NFE are highlighted. means using LLM rewritten prompts for GenEval. Qwen-Image-TwinFlow in this table is under LoRA training. Qwen-Image-Lightning generates almost identical images for the same prompt when evaluating on GenEval and DPG-Bench.
Method NFE $\downarrow$ Image Generation
GenEval $\uparrow$ DPG-Bench $\uparrow$ WISE $\uparrow$
Chameleon (Team, 2024) - 0.39 - -
[…]
Janus-Pro (Chen et al., 2025d) - 0.80 84.19 0.35
[…]
OpenUni-L-512 (Wu et al., 2025c) $\times2 0$.85 81.54 0.52
[…]
Qwen-Image-Lightning (ModelTC, 2025) 1 0.85 87.79 0.51
OpenUni-RCGM-512 (Sun & Lin, 2025) 2 0.85 80.15 0.50
OpenUni-RCGM-512 (Sun & Lin, 2025) 1 0.80 76.40 0.45
2 0.85 79.82 0.50
OpenUni-TwinFlow-512 (Ours) 1 0.83 79.07 0.48
Qwen-Image-RCGM (Sun & Lin, 2025) 2 0.82 84.09 0.50
Qwen-Image-RCGM (Sun & Lin, 2025) 1 0.52 59.50 0.30
2 0.87 / 0.91 87.64 0.57
Qwen-Image-TwinFlow (Ours) 1 0.86 / 0.90 86.52 0.54
Further exploration on 20B full-parameter training on Qwen-Image.
Tab. 3 demonstrates the scalability and performance advantages of TwinFlow on the large-scale Qwen-Image-20B. Existing distribution matching methods, such as VSD (Wang et al., 2023), DMD (Yin et al., 2024b), and SiD (Zhou et al., 2024), typically require maintaining three separate model copies (generator, real score, and fake score), leading to significant memory overhead. In contrast, TwinFlow distinguishes itself through a unified design:
  1. (a)
    Simplicity and efficiency: By integrating the generator, real/fake socre estimation into a single model, TwinFlow eliminates the need for redundant parameters. This allows for full-parameter training at the 20B scale.
  2. (b)
    Performance superiority: With this unified design, TwinFlow outperforms all baselines on Qwen-Image-20B. Notably, it achieves superior generation quality with just 1-2 NFE compared to sCM (Lu & Song, 2024) and MeanFlow (Geng et al., 2025) operating at 8 NFE.
Table 3: Comparison of full-parameter training efficiency and performance between TwinFlow and baselines on text-to-image tasks using Qwen-Image 20B. The raw setting denotes that the generator, real score, and fake score are instantiated as separate models using FSDP-v2; this configuration leads to OOM. Therefore, for VSD, SiD, and DMD, the fake score estimator is implemented using LoRA ($r=64$) to ensure memory feasibility. $\star$ indicates severe diversity degradation (mode collapse), characterized by nearly identical outputs on GenEval and DPG-Bench. For sCM and MeanFlow, the Jacobian-Vector Product (JVP) is approximated via finite differences.
Method NFE $\downarrow$ Image Generation
GenEval $\uparrow$ DPG-Bench $\uparrow$ WISE $\uparrow$
[…]
VSD (Wang et al., 2023) (raw) - OOM OOM OOM
DMD (Yin et al., 2024b) (raw) - OOM OOM OOM
SiD (Zhou et al., 2024) (raw) - OOM OOM OOM
VSD (Wang et al., 2023) 1 0.67 84.44 0.22
2 0.73 86.16 0.34
DMD (Yin et al., 2024b) 1 0.81 84.31 0.47
2 0.80 84.08 0.46
SiD (Zhou et al., 2024) 1 0.77 87.05 0.42
2 0.78 86.94 0.41
sCM (Lu & Song, 2024) (JVP-free) 4 0.62 85.37 0.44
8 0.60 85.54 0.45
MeanFlow (Geng et al., 2025) (JVP-free) 4 0.44 83.28 0.34
8 0.49 83.81 0.37
Ours 1 0.85 85.44 0.51
2 0.86 86.35 0.55
Ours (longer training) 1 0.89 87.54 0.57
2 0.90 87.80 0.59
Figure 3: Visualization of images generated by Qwen-Image and Qwem-Image-TwinFlow w.r.t. NFEs. Qwen-Image-TwinFlow is capable of generating high-quality images with just 1 NFE, which is better than the original Qwen-Image’s performance at 16 NFEs. Furthermore, when comparing 2-NFE results to the 32-NFE outputs of Qwen-Image, our method demonstrates better visual details. See prompts in App. D.3.
Discussion on open-source community efforts.
To the best of our knowledge, Qwen-Image-Lightning (ModelTC, 2025) is the only open-source few-step model on large models. It is developed using DMD2 (Yin et al., 2024a) but removing GAN loss. This also indirectly reflects the high cost associated with using GAN loss. However, we observe that Qwen-Image-Lightning suffers from severe mode collapse: when given the same prompt but different noise inputs, the generated images remain nearly identical across runs. This lack of diversity is empirically demonstrated in the visual comparisons provided in App. D.1.
Exploration on image editing.
Due to resource constraints, we conducted a preliminary exploration of our TwinFlow’s capabilities in image editing using a small tuning dataset of approximately 15K editing pairs. Despite the limited scale, our results (see Tab. 8) demonstrate that TwinFlow can convert Qwen-Image-Edit (Wu et al., 2025a) into a 4-NFE editing model. This suggests that with access to more diverse editing datasets, we anticipate substantial further improvements in both fidelity and versatility of edited outputs.
4.3 Image Generation on Dedicated Text-to-image Models
To validate our method’s versatility, we also benchmark it on traditional text-to-image generation. As shown in Tab. 4, we first benchmark against pretrained multi-step models (typically requiring >40-NFE). Following the categorization in Tab. 1, we compare against SOTA few-step models, grouped by their reliance on auxiliary components: those trained with versus without auxiliary models. Critically, full-parameter tuning on SANA-0.6B/1.6B backbones enables high-fidelity image generation in just 1-2 NFE.
Table 4: System-level comparison of TwinFlow with text-to-image models in efficiency and performance. Throughput (batch=10) and latency (batch=1) are benchmarked on a single A100 with BF16 precision. The best and second best results across 1-NFE are highlighted. means results tested by ourselves.
Method NFE $\downarrow$ Throughput $\uparrow$ Latency (s) $\downarrow$ #Params GenEval $\uparrow$ DPG-Bench $\uparrow$
(samples/s)
Pretrained multi-step models
[…]
PixArt-$\Sigma$ (Chen et al., 2024a) $\times2 0$.40 2.7 0.6B 0.54 80.5
[…]
Playground v3 (Liu et al., 2024) - 0.06 15.0 24B 0.76 87.0
[…]
SANA-1.5 (Xie et al., 2025a) $\times2 0$.26 4.2 4.8B 0.81 84.7
[…]
Few-step models (training w/ auxiliary models)
SDXL-DMD2 (Yin et al., 2024a) 2 2.89 0.40 0.9B 0.58 -
FLUX-Schnell (Labs, 2024) 2 0.92 1.15 12.0B 0.71 -
SANA-Sprint-0.6B (Chen et al., 2025b) 2 6.46 0.25 0.6B 0.76 81.5
SANA-Sprint-1.6B (Chen et al., 2025b) 2 5.68 0.24 1.6B 0.77 82.1
PixArt-DMD (Chen et al., 2024a) 1 4.26 0.25 0.6B 0.45 -
SDXL-DMD2 (Yin et al., 2024a) 1 3.36 0.32 0.9B 0.59 -
FLUX-Schnell (Labs, 2024) 1 1.58 0.68 12.0B 0.69 -
SANA-Sprint-0.6B (Chen et al., 2025b) 1 7.22 0.21 0.6B 0.72 78.6
SANA-Sprint-1.6B (Chen et al., 2025b) 1 6.71 0.21 1.6B 0.76 80.1
Few-step models (training w/o auxiliary models)
SDXL-LCM (Luo et al., 2023) 2 2.89 0.40 0.9B 0.44 -
PixArt-LCM (Chen et al., 2024b) 2 3.52 0.31 0.6B 0.42 -
PCM (Wang et al., 2024a) 2 2.62 0.56 0.9B 0.55 -
SD3.5-Turbo (Esser et al., 2024a) 2 1.61 0.68 8.0B 0.53 -
RCGM-0.6B (Sun & Lin, 2025) 2 6.50 0.26 0.6B 0.85 80.3
RCGM-1.6B (Sun & Lin, 2025) 2 5.71 0.25 1.6B 0.84 79.1
TwinFlow-0.6B (Ours) 2 6.50 0.26 0.6B 0.84 79.7
TwinFlow-1.6B (Ours) 2 5.71 0.25 1.6B 0.83 79.6
SDXL-LCM (Luo et al., 2023) 1 3.36 0.32 0.9B 0.28 -
PixArt-LCM (Chen et al., 2024b) 1 4.26 0.25 0.6B 0.41 -
PCM (Wang et al., 2024a) 1 3.16 0.40 0.9B 0.42 -
SD3.5-Turbo (Esser et al., 2024a) 1 2.48 0.45 8.0B 0.51 -
RCGM-0.6B (Sun & Lin, 2025) 1 7.30 0.23 0.6B 0.80 77.2
RCGM-1.6B (Sun & Lin, 2025) 1 6.75 0.22 1.6B 0.78 76.5
TiM (Wang et al., 2025) 1 - - 0.8B 0.67 75.0
TwinFlow-0.6B (Ours) 1 7.30 0.23 0.6B 0.83 78.9
TwinFlow-1.6B (Ours) 1 6.75 0.22 1.6B 0.81 79.1
  1. (a)
    1-NFE setting: The efficacy of our method is particularly pronounced in the more demanding 1-NFE inference setting. Here, our models (0.6B: 0.83, 1.6B: 0.81 on GenEval) significantly outperform other leading 1-NFE methods, such as SANA-RCGM (0.78) (Sun & Lin, 2025), SANA-Sprint (0.76) (Chen et al., 2025b), FLUX-Schnell (0.69) (Labs, 2024), and SDXL-DMD2 (0.59) (Yin et al., 2024a). Notably, our 1-NFE TwinFlow-0.6B (GenEval: 0.83) exceeds the generation quality of the 40-NFE SANA-1.5-4.8B (Xie et al., 2025a) model while offering substantially greater computational efficiency.
  2. (b)
    2-NFE setting: In the 2-NFE configuration, TwinFlow-0.6B achieves a throughput of 6.50 samples/s and a latency of 0.26s, performance metrics comparable to the originally reported SANA values. On the GenEval benchmark, our model attains a score of 0.84, surpassing not only the SANA-Sprint series (0.76 and 0.77) but also powerful multi-step models like SANA-1.5 (0.81) and Playground v3 (0.76). Our models also demonstrate competitive performance on DPG-Bench, with scores of 79.7 for the 0.6B variant and 79.6 for the 1.6B variant.
Our TwinFlow-0.6B/1.6B achieves state-of-the-art text-to-image generation performance on the GenEval benchmark using just 1-NFE, surpassing both SANA-Sprint and RCGM. While we slightly underperform on DPG-Bench relative to SANA-Sprint, due to SANA-Sprint’s reliance on extensive, proprietary training data. We believe this gap is primarily data-driven and can be effectively closed by training on larger, higher-quality datasets.
4.4 Ablation Study and Analysis
Influence of $\lambda$. As described in Sec. 3.3, $\lambda$ is designed to control the sample distribution of $\mathcal{L}_{\mathrm{base}}$ and $\mathcal{L}_{\mathrm{TwinFlow}}$. In Fig. 4(a), we visualize the DPG-Bench performance w.r.t. $\lambda$ at 1-NFE and 2-NFE. We observed that as $\lambda$ increases from 0, the performance on DPG-Bench initially increases and then decreases, reaching its peak at approximately $\lambda=\nicefrac{{1}}{{3}}$. These results indicate that appropriately balancing samples in the local batch helps improve the model performance.
Impact of $\mathcal{L}_{\mathrm{TwinFlow}}$ on different models. We conduct an ablation study to analyze the impact on text-to-image performance of using $\mathcal{L}_{\mathrm{TwinFlow}}$ on different models. As illustrated in Fig. 4(b), incorporating $\mathcal{L}_{\mathrm{TwinFlow}}$ significantly enhances performance: it improves 1-NFE performance for the text-to-image task across OpenUni, SANA, and especially Qwen-Image (from 59.50 to 86.52).
Effect of training steps vs. NFE. As illustrated in Fig. 4(c), the experimental results demonstrate that as the number of training steps increases, the “comfort regime” for optimal sampling steps shifts accordingly. Notably, performance on GenEval improvements are observed across both 1-step and few-step sampling scenarios, with significant gains achieved as training progresses, which shows the effectiveness of $\mathcal{L}_{\mathrm{TwinFlow}}$.
(a) Influence of $\lambda$.
(b) Impact of $\mathcal{L}_{\mathrm{TwinFlow}}$.
(c) Training steps vs. NFE.
Figure 4: Ablation studies of TwinFlow. Ablation presented in (a) and (c) are conducted on Qwen-Image-TwinFlow. Results shown in (b) are trained on the same dataset but with different models.
5 Conclusion and Limitations
In this paper, we introduce TwinFlow, a simple yet effective framework for training large-scale few-step continuous generative models. Our method stands out for its high simplicity compared to other few-step approaches, such as the DMD-series, as it eliminates the need for auxiliary trained components like GAN discriminators or frozen teacher models. This design allows for straightforward 1-step or few-step training on large models, making it particularly accessible and efficient. Through extensive experiments across different scales and tasks, we demonstrate that TwinFlow delivers high-quality generation capabilities in text-to-image synthesis on large models. Despite these promising results, several limitations remain to be addressed. First, the scalability of TwinFlow to more diverse tasks, such as image editing, has not been effectively explored. Second, its adaptability to more diverse modalities, including video and audio generation, requires further validation. Addressing these challenges could significantly enhance the applicability and performance of TwinFlow in broader contexts, paving the way for more robust and versatile generative models.
Appendix A Related Work
Multi-step generative methods.
Diffusion (Ho et al., 2020; Dhariwal & Nichol, 2021) and flow matching (Lipman et al., 2022) models have shown impressive performance in image generation. They progressively transport a simple noise distribution to the data distribution, either by reversing a noising process or integrating a probability-flow ODE. However, this iterative nature introduces bottleneck in inference efficiency, as generating an image requires numerous sequential evaluation steps.
Few-step generative methods.
Various methods tend to accelerate the sampling process, such as diffusion distillation (Luhman & Luhman, 2021; Salimans & Ho, 2022) and consistency distillation (Song et al., 2023; Song & Dhariwal, 2023). They typically train a student model to approximate the ODE sampling trajectory of a frozen teacher model in fewer sampling steps. While effective at moderate NFEs, these approaches depend on a frozen teacher, and quality often drops sharply in the extreme few-step regime (< 4 NFEs). Incorporating GAN-like loss into distillation (e.g., CTM (Kim et al., 2023), ADD/LADD (Sauer et al., 2024b; a), DMD/DMD2 (Yin et al., 2024b; a)) can improve sharpness and alignment at few steps. Yet these frameworks increase training complexity and instability: they typically introduce auxiliary modules (discriminators, fake-sample score functions) and still rely on a frozen teacher, leading to higher memory overhead and sensitivity to hyperparameters. For ultra-large models, this added complexity often translates to out-of-memory failures or brittle training dynamics.
Few-step applications in large generative models.
The tension between speed and quality is amplified in large-scale systems. For instance, Qwen-Image-20B (Wu et al., 2025a) typically requires 100 NFEs, leading to substantial latency ($\sim$40s on a single A100 for $\times1024$ resolution). Recent works distill to cut NFEs while preserving compositionality and aesthetics: LCM-style distillation for latent models (Luo et al., 2023), large-scale text-to-image distillation pipelines (e.g., PixArt-Delta (Chen et al., 2024b), SDXL distillation (Lin et al., 2024), FLUX-schnell (Labs, 2024)), and hybrid frameworks such as SANA-Sprint (Chen et al., 2025b) that combine teacher guidance with adversarial signals. Recently, Hunyuan-Image-2.1 (Team, 2025) explore MeanFlow (Geng et al., 2025) for mid-step acceleration (16 NFEs). Nevertheless, when targeting 1-2 NFEs on $\times1024$ text-to-image with 10B-20B backbones, these pipelines face practical barriers: dependence on frozen teachers, extra discriminators or score networks, unstable adversarial training, and prohibitive memory costs that hinder straightforward scaling.
Appendix B Detailed Experiments
B.1 Detailed Implementation on Multimodal Generative Models
Training Datasets.
For training OpenUni (Wu et al., 2025c) and Qwen-Image (Wu et al., 2025a), we used the same datasets as in our text-to-image experiments but excluded ShareGPT-4o-Image (Chen et al., 2025c). For exploration on full-parameter training on Qwen-Image, since DMD (Yin et al., 2024b) employs a regression loss, we need to pre-generate offline data using Qwen-Image. To ensure a fair comparison with established baselines, we randomly sample 50K prompts from the datasets we used, and generate images using Qwen-Image. For Qwen-Image-Edit (Wu et al., 2025a), we used only the part of editing split of the ShareGPT-4o-Image dataset.
Training configurations.
We performed full-parameter fine-tuning experiments on OpenUni, using an image resolution of $\times512$. With a batch size of 128, the model was trained for 600,000 steps. For other training configurations, please refer to Tab. 6.
For LoRA fine-tuning on Qwen-Image and Qwen-Image-Edit. We set the rank ($r$) and alpha ($\alpha$) to 64 for Qwen-Image, and to 64 and 32, respectively, for Qwen-Image-Edit. This LoRA setup comprises approximately 420M trainable parameters. A comprehensive list of training configuration is provided in Tab. 6. For full-parameter training on Qwen-Image, we add a additional time embedder to serve as the target timestep condition.
Additional detailed results on GenEval.
As detailed in Tab. 5, our Qwen-Image-TwinFlow achieves 0.86 with 1-NFE. Notably, when using LLM rewritten prompts, our GenEval score comes to 0.90, which is very close to Qwen-Image-RL (Wu et al., 2025a) (0.91). Significant score increases are observed in the Colors and Attribute Binding subtasks. This indicates that our model exhibits enhanced image generation capabilities when processing long input instructions.
Table 5: Detailed evaluation results on GenEval for text-to-image models. Qwen-Image-Lightning are evaluated with 1-NFE. means using LLM rewritten prompts for GenEval.
[…]
Object Object Binding
SEED-X (Ge et al., 2024) 0.97 0.58 0.26 0.80 0.19 0.14 0.49
Emu3-Gen (Wang et al., 2024b) 0.98 0.71 0.34 0.81 0.17 0.21 0.54
JanusFlow (Ma et al., 2025) 0.97 0.59 0.45 0.83 0.53 0.42 0.63
Show-o (Xie et al., 2024b) 0.99 0.80 0.66 0.84 0.31 0.50 0.68
OmniGen (Xiao et al., 2024) 0.98 0.84 0.66 0.74 0.40 0.43 0.68
Janus-Pro-7B (Chen et al., 2025d) 0.99 0.89 0.59 0.90 0.79 0.66 0.80
OpenUni-512 (Wu et al., 2025c) 0.99 0.91 0.77 0.90 0.75 0.76 0.85
Bagel (Deng et al., 2025) 0.99 0.94 0.81 0.88 0.64 0.63 0.82
OmniGen2 (Wu et al., 2025b) 1.00 0.95 0.64 0.88 0.55 0.76 0.80
Show-o2-7B (Xie et al., 2025b) 1.00 0.87 0.58 0.92 0.52 0.62 0.76
Qwen-Image (Wu et al., 2025a) 0.99 0.92 0.89 0.88 0.76 0.77 0.87
Qwen-Image-Lightning (ModelTC, 2025) 0.99 0.89 0.85 0.87 0.75 0.76 0.85
OpenUni-TwinFlow-512 (1-NFE) 0.99 0.91 0.69 0.90 0.79 0.72 0.83
Qwen-Image-TwinFlow (1-NFE) 1.00 0.91 0.84 0.90 0.75 0.74 0.86
Qwen-Image-TwinFlow (1-NFE) 0.99 0.94 0.87 0.96 0.78 0.83 0.90
B.2 Detailed Implementation on Dedicated Text-to-image Models
Table 6: Detailed training configurations for experiments on text-to-image models and multimodal generative models.
Configuration SANA-0.6B SANA-1.6B OpenUni-512 Qwen-Image (LoRA) Qwen-Image (Full) Qwen-Image-Edit
Optimizer Settings
Optimizer RAdam RAdam AdamW AdamW AdamW AdamW
Learning Rate $\times 10^{-4} 1×1$$\times 10^{-4}$ $\times 10^{-4} 1×1$$\times 10^{-4}$ $\times 10^{-5} 1×1$$\times 10^{-4}$
Weight Decay 0 0 0 0 0 0
$(\beta_{1},\beta_{2})$ (0.9, 0.95) (0.9, 0.99) (0.9, 0.95) (0.9, 0.99) (0.9, 0.99) (0.9, 0.99)
Training Details
Batch Size 128 128 128 64 32 24
Training Steps 30000 30000 60000 7000 3000, 10000 (longer training) 7000
Learning Rate Scheduler Constant Constant Constant Constant Constant Constant
Gradient Clipping - - - 1.0 1.0 1.0
Random Seed 42 42 42 42 42 42
LoRA $r$ - - - 64 - 64
LoRA $\alpha$ - - - 64 - 32
EMA Decay Rate 0.99 0.99 0.99 0 0.99 0
Training Datasets.
Considering training datasets, we use BLIP-3o-60K (Chen et al., 2025a), Echo-4o (w/o multi-reference split) (Ye et al., 2025), and ShareGPT-4o-Image (Chen et al., 2025c). Together, these three instruction tuning datasets comprise approximately 200,000 text-to-image samples.
Training configurations.
All experiments on SANA-0.6B/1.6B backbones are conducted with full-parameter tuning. we fine-tuned these two models for 30,000 steps, using batch sizes of 128 and 64 respectively. Other detailed training configurations are provided in Tab. 6.
Additional detailed results on GenEval.
We list the detailed results of TwinFlow-0.6B/1.6B on GenEval (Ghosh et al., 2023) along with other SOTA multi-step text-to-image models (except FLUX-Schnell) in Tab. 7. It demonstrates except the overall score, our TwinFlow-0.6B/1.6B also outperforms these multi-step models with 1-NFE in sub tasks such as Position (0.6B: 0.84, 1.6B: 0.79) and Attribute Binding (0.6B: 0.70, 1.6B: 0.68).
Table 7: Detailed evaluation results on GenEval for text-to-image models.
[…]
Object Object Binding
SDXL (Lin et al., 2024) 0.98 0.74 0.39 0.85 0.15 0.23 0.55
PixArt-$\Sigma$ (Chen et al., 2024a) 0.98 0.59 0.50 0.80 0.10 0.15 0.52
SD3-Medium (Esser et al., 2024a) 0.98 0.74 0.63 0.67 0.34 0.36 0.62
FLUX-Dev (Labs, 2024) 0.98 0.81 0.74 0.79 0.22 0.45 0.66
FLUX-Schnell (Labs, 2024) 0.99 0.92 0.73 0.78 0.28 0.54 0.71
SD3.5-Large (Esser et al., 2024a) 0.98 0.89 0.73 0.83 0.34 0.47 0.71
Lumina-Image-2.0 (Qin et al., 2025) - 0.87 0.67 - - 0.62 0.73
SANA-0.6B (Xie et al., 2024a) 0.99 0.76 0.64 0.88 0.18 0.39 0.64
SANA-1.6B (Xie et al., 2024a) 0.99 0.77 0.62 0.88 0.21 0.47 0.66
TwinFlow-0.6B (1-NFE) 0.98 0.90 0.68 0.89 0.84 0.70 0.83
TwinFlow-1.6B (1-NFE) 0.99 0.88 0.65 0.86 0.79 0.68 0.81
B.3 Exploration on Image Editing
We conducted a preliminary exploration of our method on image editing tasks using a subset of approximately 15,000 square images from Chen et al. (2025c). We fine-tuned the Qwen-Image-Edit model (Wu et al., 2025a) using LoRA at a fixed $\times512$ resolution, with training configurations detailed in Table 6. Note that we use a low and fixed resolution during training; during testing, we use a resolution that is the same as the input image size. Our Qwen-Image-Edit-TwinFlow can effectively edit images in 2 to 4 NFEs. With 2 NFEs, it achieves a score of 3.47 on the ImgEdit, surpassing all multi-step models except Qwen-Image-Edit itself.
Despite the naive usage of the dataset and training strategy, these results highlight the significant potential of our approach. We believe that with further scaling, our method can achieve strong performance in a single-step (1-NFE) setting on image editing tasks.
Table 8: Comparison of TwinFlow with unified multimodal models in performance on image editing tasks. In GEdit-Bench, G_SC measures Semantic Consistency, G_PQ evaluates Perceptual Quality, and G_O reflects the Overall Score. All metrics are evaluated by GPT-4.1.
Method NFE $\downarrow$ Image Editing
GEdit-EN (Full set) $\uparrow$ ImgEdit $\uparrow$
G_SC G_PQ G_O
[…]
Qwen-Image-Edit-TwinFlow 4 5.95 6.97 5.91 3.55
Qwen-Image-Edit-TwinFlow 2 5.94 6.81 5.85 3.47
Appendix C Theoretical Analysis
C.1 Transformation from Score to Velocity
In this section, we derive the equation between score and velocity. According to Equation 2 in Sun et al. (2025) with flow matching, we have:
$\boldsymbol{f}^{\mathbf{x}}(\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t),\mathbf{x}_{t},t)$ $=\frac{\alpha(t)\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)-\hat{\alpha}(t)\cdot\mathbf{x}_{t}}{\alpha(t)\cdot\hat{\gamma}(t)-\hat{\alpha}(t)\cdot\gamma(t)}=\frac{t\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)-\mathbf{x}_{t}}{(t\cdot(-1)-1\cdot(1-t))}$
$=\frac{t\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)-\mathbf{x}_{t}}{(-1)}=\mathbf{x}_{t}-t\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)\,.$ (11)
According to Theorem 2 in Sun et al. (2025), we have:
$f^{\mathbf{x}}(\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t),\mathbf{x}_{t},t)$ $=\frac{\mathbf{x}_{t}+\alpha^{2}(t)\,\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})}{\gamma(t)}=\frac{\mathbf{x}_{t}+t^{2}\,\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})}{(1-t)}=\frac{\mathbf{x}_{t}+t^{2}\,\mathbf{s}(\mathbf{x}_{t})}{1-t}\,$, (12)
where $\mathbf{s}(\mathbf{x}_{t})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$. By simple transposition of $\mathbf{s}(\mathbf{x}_{t})$ term, we can get:
$\frac{\mathbf{x}_{t}+t^{2}\,\mathbf{s}(\mathbf{x}_{t})}{1-t}$ $=\mathbf{x}_{t}-t\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$
$\implies t^{2}\mathbf{s}(\mathbf{x}_{t})$ $=(1-t)\left(\mathbf{x}_{t}-t\cdot\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)\right)-\mathbf{x}_{t}=-t\mathbf{x}_{t}+(t^{2}-t)\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$
$\implies\mathbf{s}(\mathbf{x}_{t})$ $=\frac{-t\mathbf{x}_{t}+(t^{2}-t)\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)}{t^{2}}=\boxed{-\frac{\mathbf{x}_{t}+(1-t)\boldsymbol{F}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)}{t}}\,.$ (13)
Appendix D Visualization Results
D.1 Comparison of Qwen-Image-TwinFlow and Qwen-Image-Lightning
As Fig. 5 shown, our comparative analysis reveals a notable limitation in Qwen-Image-Lightning’s generation. The model produces images with very low diversity; outputs are often highly similar and visually repetitive even when initialized with different latent noise. Our Qwen-Image-TwinFlow does not exhibit model collapse, demonstrating the ability to generate a rich variety of high-quality images. We also quantify the similarity by using LPIPS metric in Tab. 9. We perform the diversity evaluation on GenEval. Specifically, for each prompt, we compute the average pairwise LPIPS distance among the 4 generated samples. The final score is the mean of all scores across the entire samples. The results quantify that Qwen-Image-Lightning shows obvious diversity degradation.
Table 9: Quantification of diversity using LPIPS score for Qwen-Image-Lightning and Qwen-Image-TwinFlow. Qwen-Image-Lightning shows obvious diversity degradation.
Model NFE LPIPS $\uparrow$
Qwen-Image-Lightning 1 0.2996
Qwen-Image-TwinFlow 1 0.5044
Qwen-Image-Lightning 2 0.3046
Qwen-Image-TwinFlow 2 0.5188
Figure 5: Comparison between Qwen-Image-TwinFlow and Qwen-Image-Lightning (1-NFE). The prompts and generated images are sourced from DPG-Bench. We observe that Qwen-Image-Lightning tend to generate very similar images though noise is different, which hurts diversity. Our model remains diversity and high quality generation.
D.2 Visualization Results across Training Steps
Figure 6: Visualizations of 1-NFE images generated by Qwen-Image-TwinFlow w.r.t. training steps. In the early stages of training, our method converges rapidly, and the generated images begin to take shape (200 to 400 steps); as training progresses, our method gradually optimize the visual details (800 to 6400 steps).
As illustrated in Fig. 6, our method exhibits a two-stage training dynamic. Initially (200 to 400 steps), it demonstrates rapid convergence in 1-NFE performance, quickly establishing a strong baseline. Subsequently (800 to 6400 steps), the training process shifts towards refining finer visual details, leading to a steady enhancement of image fidelity. This highlights our approach’s efficiency in achieving strong initial results and its capacity for continued improvement with extended training.
D.3 Selected Prompts Used for Visualization
The prompts used to generate the results shown in Fig. 1 and Fig. 3 are detailed in this section to ensure reproducibility.
1A cinematic vertical composition of a bustling street in Central Hong Kong, featuring vibrant red taxis driving along the road. The scene is framed by towering modern skyscrapers with reflective glass facades, creating an urban jungle atmosphere. The cityscape glows with neon signs and soft ambient light, capturing the essence of Hong Kongs iconic night-time energy. On the road surface, painted traffic markings texts: SLOW’. The sky above has a gradient transitioning from deep twilight blue to warm orange hues near the horizon, adding depth and drama to the image. Rendered in hyper-realistic style with rich colors, intricate textures, and high contrast lighting for maximum impact.
2Classic Baroque-style still life painting, a woven wicker basket overflowing with fresh fruits including red and green grapes, ripe apples, plums, quinces, and a yellow pear, adorned with grape leaves and vines. The basket sits on a draped stone or wooden table covered with a dark blue cloth, with scattered fruits and berries around it. Rich, dramatic lighting highlights the textures and colors of the fruit, creating deep shadows and soft highlights. A luxurious red curtain drapes in the background, adding depth and contrast. Realistic, highly detailed, oil painting style, reminiscent of 17th-century Dutch or Flemish masters such as Jan Davidsz de Heem or Caravaggio. Warm, earthy tones, meticulous attention to detail, and a sense of abundance and natural beauty. Ultra HD, 4K, cinematic composition.
3Clean white brick wall, vibrant colorful spray-paint graffiti covering entire surface: top giant bubble letters ’1 Step Generation’, below stacked TwinFlow’, Made Easy in rainbow palette, fresh wet paint drips, daylight urban photography, realistic light and shadow. Ultra HD, 4K, cinematic composition.
4A close-up realistic selfie of three cats of different breeds in front of the iconic Big Ben, each with a different expression, taken during the blue hour with cinematic lighting. The animals are close to the camera, heads touching, mimicking a selfie pose, displaying joyful, surprised, and calm expressions. The background showcases the complete architectural details of the [landmark], with soft light and a warm atmosphere. Shot in a photorealistic, cartoon style with high detail.
5Starbucks miniature diorama shop. The roof is made of oversized coffee beans, and above the windows is a huge Starbucks sign. A vendor is handing coffee to customers, and the ground is covered with many coffee beans. Handmade polymer clay sculpture, studio macro photograph, soft lighting, shallow depth of field. Ultra HD, 4K, cinematic composition.
6A whimsical scene featuring a capybara joyfully riding a sleek, modern rocket. The capybara is holding a sign with both hands, the text on the sign boldly and eye-catchingly reading QWEN-IMAGE-20B’. The capybara looks thrilled, sporting a playful grin as it soars through a vibrant sky filled with soft, pastel clouds and twinkling stars. The rocket leaves a trail of sparkling, colorful smoke behind it, adding to the magical atmosphere. Ultra HD, 4K, cinematic composition.A still frame from a black and white movie, featuring a man in classic attire, dramatic high contrast lighting, deep shadows, retro film grain, and a nostalgic cinematic mood. Ultra HD, 4K, cinematic composition.
7Close-up portrait of a young woman with light skin and long brown hair, looking directly at the camera. Her face is illuminated by dramatic, slatted sunlight casting shadows across her features, creating a pattern of light and shadow. Her eyes are a striking green, and her lips are slightly parted, with a natural pink hue. The background is a soft, dark gradient, enhancing the focus on her face. The lighting is warm and golden. Ultra HD, 4K, cinematic composition.
8A field of vibrant red poppies with green stems under a blue sky.
9A small blue-gray butterfly with black stripes rests on a white and yellow flower against a blurred green background.
10A grey tabby cat with yellow eyes rests on a weathered wooden log under bright sunlight
D.4 High Resolution Visualization
In this section, we showcase further qualitative results from Qwen-Image-TwinFlow to highlight its generative capabilities. To ensure an unbiased representation, the generation prompts were chosen at random, and the resulting visualizations are presented without curation or cherry-picking.
Figure 7: Visualization of Qwen-Image-TwinFlow (NFE=4). Each image is of $\times1328$ resolution.
Figure 8: Visualization of Qwen-Image-TwinFlow (NFE=4). Each image is of $\times1328$ resolution.
Figure 9: Visualization of Qwen-Image-TwinFlow (NFE=4). Each image is of $\times1328$ resolution.
Figure 10: Visualization of Qwen-Image-TwinFlow (NFE=4). Each image is of $\times1328$ resolution.
Figure 11: Visualization of Qwen-Image-TwinFlow (NFE=4). Each image is of $\times1328$ resolution.
D.5 Fake Trajectory Visualization
Figure 12: Visualization of the fake trajectory (NFE=20). The fake images have significant visual difference comparing to real images.